OpenStudy (anonymous):

Evaluate 12÷(-3)(2) + 5 is it -13?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

use BODMAS

OpenStudy (anonymous):

lol,no. 1st 12/(-3) = -4.,then (-4)*2=-8 and last -8 +5= -3

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

No Lalaly,multiplication and division are going at same time. When we have both of them we are going from left to right.